# Legendrian Lavrentiev links

### v1

@@ 2. Lavrentiev curves
The notion of a Lavrentiev curve can be found in [17, Theorem 8 2)]. Definition 2.29 in [18] of a quasiconvex metric space generalizes the concept of a Lavrentiev curve.
@@ 3.2. Pushing the regular projection smoothly to correct the integral
Let $t_{-}$ be some non-negative real number. Then by Stokes theorem (see [25, Chapter IX, Theorems 5A and 7A])
@@ 4. Smoothing Lavrentiev curves on surfaces
Now we address condition (5). First, we orient the surface $S$ using the 2-form $d\beta$ which is nonzero everywhere by the assumption of the present proposition. Second, we note that the isotopy $F_{0}$ always pushes the subarc $L_{0}$ to one side of the curve. So by Stokes theorem (see [25, Chapter IX, Theorems 5A and 7A]) the integral $\int\limits_{F(L\times\{t\})}\beta$ is a monotone function on $t$. We will smoothly push the curve to the opposite side to make this integral constant.
@@ 5.1. Case $r=+\infty$
By Stokes theorem (see [25, Chapter IX, Theorems 5A and 7A])
@@ 8. Contactomorphisms
In this section we assume the ambient manifold to be compact and connected if not stated otherwise. We continue to use Convention 2.9.
We equip the set of all bi-Lipschitz self-homeomorphisms of the manifold $M$ with the metric
$\mathrm{dist}(f,g)=\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p))+\sup\limits_{p\neq q}\left|\ln\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(g(p),g(q))}\right|.$
@@
The metric space of bi-Lipschitz self-homeomorphisms is complete.
@@
Pick a Cauchy sequence $f_{n}.$
It is also a Cauchy sequence for the $C^{0}$-part of our metric. So it $C^{0}$-converges to a surjective continuous map $f.$ If $f$ is not bi-Lipschitz, for any $C$ there exists a pair of points $p,q$ such that $\left|\ln\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(p,q)}\right|>C.$ Since the sequence $f_{n}$ $C^{0}$-converges to $f$, $\left|\ln\frac{\mathrm{dist}(f_{n}(p),f_{n}(q))}{\mathrm{dist}(p,q)}\right|>C$ for sufficiently large $n$. Making $C$ arbitrarily large we see that the sequence $f_{n}$ is unbounded, which is a contradiction. Thus $f$ is a bi-Lipschitz homeomorphism.
It remains to prove that $\mathrm{dist}(f_{n},f)\to 0.$ The $C^{0}$-part of this distance is small by construction. Take any small $\varepsilon>0.$ There exists $N$ such that for any $n,m$ greater than $N$ we have $\mathrm{dist}(f_{n},f_{m})<\varepsilon.$ Take any two distinct points $p,q$ and a number $n$ greater than $N$. There exists $m$ greater than $N$ such that $\mathrm{dist}(f_{m}(p),f(p))<\varepsilon\cdot\mathrm{dist}(f(p),f(q))$ and $\mathrm{dist}(f_{m}(q),f(q))<\varepsilon\cdot\mathrm{dist}(f(p),f(q)).$ Therefore
(19) $\ln\left|\frac{\mathrm{dist}(f_{n}(p),f_{n}(q))}{\mathrm{dist}(f(p),f(q))}\right|\leq\ln\left|\frac{\mathrm{dist}(f_{n}(p),f_{n}(q))}{\mathrm{dist}(f_{m}(p),f_{m}(q))}\right|+\ln\left|\frac{\mathrm{dist}(f_{m}(p),f_{m}(q))}{\mathrm{dist}(f(p),f(q))}\right|\leq\\ \leq\ln\left|\frac{\mathrm{dist}(f_{n}(p),f_{n}(q))}{\mathrm{dist}(f_{m}(p),f_{m}(q))}\right|-\ln(1-2\varepsilon)<\varepsilon-\ln(1-2\varepsilon).$
Lemma 8.2.
The space of $C^{1}$-diffeomorphisms is a closed subspace of the space of bi-Lipschitz homeomorphisms. The restriction of the metric to this subspace equals
$\mathrm{dist}(f,g)=\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p))+\sup\limits_{p\in M}\left(\left|\ln||D_{p}(f\circ g^{-1})||\right|,\left|\ln||D_{p}(g\circ f^{-1})||\right|\right).$
@@
It is known that any diffeomorphism is bi-Lipschitz. Let us prove the formula for the restriction.
Let $f$, $g$ be two $C^{1}$-diffeomorphisms.
$\mathrm{dist}(f,g)=\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p))+\sup\limits_{p\neq q}\left|\ln\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(g(p),g(q))}\right|=\\ =\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p))+\sup\limits_{p\neq q}\left|\ln\frac{\mathrm{dist}(f\circ g^{-1}(p),f\circ g^{-1}(q))}{\mathrm{dist}(p,q)}\right|.$
Let $\gamma(s)$ be a smooth arc realizing the distance between two distinct points $p$ and $q$. Then
[…]
If there is no such arc $\gamma$ we can obtain the same bound by choosing a sequence of smooth arcs connecting the points $p$ and $q$ whose length tends to the distance between $p$ and $q.$ By the same argument for the inverse map $g\circ f^{-1}$
$\frac{\mathrm{dist}\left(f\circ g^{-1}(p),f\circ g^{-1}(q)\right)}{\mathrm{dist}(p,q)}\geq\sup\limits_{x\in M}||D_{x}\left(g\circ f^{-1}\right)||^{-1}.$
So
$\sup\limits_{p\neq q}\left|\ln\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(g(p),g(q))}\right|\leq\sup\limits_{x\in M}\left(\left|\ln||D_{x}(f\circ g^{-1})||\right|,\left|\ln||D_{x}(g\circ f^{-1})||\right|\right).$
Actually this inequality is an equality. To see this we can take the point $x$ and the direction $v\in T_{x}M$ where the maximum of the right side is achieved, and two points $p$ and $q$ sufficiently close to $x$ such that the points $x,p,q$ lie on a curve which is tangent to $v$.
The subspace of $C^{1}$-diffeomorphisms is closed because it is complete. ∎
Example 8.3.
Let $\rho,\varphi$ be polar coordinates on Euclidean plane. Then the homeomorphism of the unit disk $\rho\leq 1$
$(\rho,\varphi)\mapsto(\rho,\varphi+\ln\rho)$
is bi-Lipschitz and the distance from it to the subspace of $C^{1}$-diffeomorphisms is at least $\ln(1+\sqrt{2})$.
Definition 8.4.
A bi-Lipschitz self-homeomorphism of a smooth contact 3-manifold is called contact if it preserves the class of Lavrentiev Legendrian curves. We call such a homeomorphism a contactomorphism.
Recall that if $\beta$ is a 1-form on the smooth surface $S$ such that $d\beta$ is an area form, $\mathrm{pr}$ is the projection from $M=S\times\mathbb{R}$ to $S$ and $z$ is the projection $M\to\mathbb{R},$ then the kernel of the 1-form $dz+\mathrm{pr}^{*}\beta$ determines a contact structure on $M$. We equip $M$ with a product of a Riemannian metric on $S$ and Euclidean metric.
Lemma 8.5.
Let $h$ be a bi-Lipschitz homeomorphism of the smooth compact connected surface $S$ preserving the integral of the 1-form $\beta$ as above on all closed Lavrentiev curves. Let $p\in S.$ Then there is the unique contactomorphism $\widetilde{h}$ of $M=S\times\mathbb{R}$ such that $\widetilde{h}(p,z)=(h(p),z)$ for any $z\in\mathbb{R}$ and $\mathrm{pr}\circ\widetilde{h}=h\circ\mathrm{pr}.$
Proof.
Let $q\in S$ and $q\neq p.$ We connect the points $p$ and $q$ by a Lavrentiev arc $\gamma$ directed from $p$ to $q.$ By the results of Section 3.1 for any $z\in\mathbb{R}$ there exists the unique Legendrian curve having the endpoint $(q,z)$ and the projection $\gamma.$ The other endpoint of this curve is $(p,z+\int\limits_{\gamma}\beta).$ Since a contactomorphism must map Legendrian curves to Legendrian curves, we must have
The right side of this formula does not depend on $\gamma$ by our assumptions on $h$. So if we take this formula as the definition of $\widetilde{h}$ we obtain a contactomorphism. ∎
Example 8.6.
Let $\rho,\varphi,z$ be cylindrical coordinates in $\mathbb{R}^{3}.$ Consider the contact form $dz+\rho^{2}d\varphi$ on the solid cylinder $\rho\leq 1.$ Then the map
is a bi-Lipschitz contactomorphism by the previous lemma. We note that the rays $\varphi=\mathrm{const},z=0$ are mapped to Legendrian curves whose projections to the plane $z=0$ are logarithmic spirals.
@@
The space of all bi-Lipschitz contactomorphisms is closed in the space of all bi-Lipschitz homeomorphisms.
Proof.
Suppose that a sequence of bi-Lipschitz contactomorphisms $f_{n}$ converges to a bi-Lipschitz homeomorphism $f$. Let $L$ be any Lavrentiev Legendrian curve. We need to prove that $f(L)$ is Legendrian.
It is clear that $f(L)$ is a Lavrentiev curve. Since the sequence $f_{n}$ is bounded, there is a real number $C$ such that all homeomorphisms $f_{n}$ are $C$-bi-Lipschitz. We can also assume that $f$ is $C$-bi-Lipschitz.
Lemma 8.8.
@@
Let us prove that a smooth contactomorphism preserves the class of Legendrian (Lavrentiev) curves. Recall that Legendrian curves are characterized by the fact that the integral of the contact form on any subarc is zero. Since the inverse image under the contactomorphism of the contact form is the contact form multiplied by some smooth nonzero function, the integral along any Legendrian arc remains to be zero after applying the contactomorphism. ∎
@@
The ambient isotopy of a contact manifold is called contact if it is given by a continuous path in the metric space of bi-Lipschitz contactomorphisms.
Figure 9. The bi-Lipschitz homeomorphisms $(\rho,\varphi)\mapsto(\rho,\varphi+t\ln\rho)$
@@
Let us prove that the isotopy of the solid cylinder $\rho\leq 1$
$h_{t}:(\rho,\varphi,z)\mapsto\left(\rho,\varphi+t\ln\rho,z-\frac{t}{2}\rho^{2}\right),\quad t\in\mathbb{R}$
is contact. This map determines a homomorphism from $(\mathbb{R},+)$ to bi-Lipschitz contactomorphisms. So we only need to prove that $h_{t}$ is small if $t$ is small.
First, let us consider the $z$-part of this homeomorphism:
$(\rho,\varphi,z)\mapsto\left(\rho,\varphi,z-\frac{t}{2}\rho^{2}\right).$
This diffeomorphism is small because the absolute value of the function $\frac{t}{2}\rho^{2}$ is not greater than $|t|/2$ and the norm of the differential of this function is not greater than $|t|.$
So it only remains to prove that the homeomorphism of the unit disk $\rho\leq 1$
$(\rho,\varphi)\mapsto\left(\rho,\varphi+t\ln\rho\right)$
is small if $t$ is small. The details are left to the reader, see Figure 9.
@@
Since $\{h_{t}\}$ is a continuous path, it is bounded, so the contactomorphisms $h_{t}$ are bi-Lipschitz with a common constant. Therefore the curves $h_{t}(L)$ are Lavrentiev with a common constant. So $\{h_{t}(L)\}$ is a Legendrian isotopy by Proposition 3.10. ∎

### v2

@@ 2. Lavrentiev curves
The notion of a Lavrentiev curve can be found in [18, Theorem 8 2)]. Definition 2.29 in [19] of a quasiconvex metric space generalizes the concept of a Lavrentiev curve.
@@ 3.2. Pushing the regular projection smoothly to correct the integral
Let $t_{-}$ be some non-negative real number. Then by Stokes theorem (see [26, Chapter IX, Theorems 5A and 7A])
@@ 4. Smoothing Lavrentiev curves on surfaces
Now we address condition (5). First, we orient the surface $S$ using the 2-form $d\beta$ which is nonzero everywhere by the assumption of the present proposition. Second, we note that the isotopy $F_{0}$ always pushes the subarc $L_{0}$ to one side of the curve. So by Stokes theorem (see [26, Chapter IX, Theorems 5A and 7A]) the integral $\int\limits_{F(L\times\{t\})}\beta$ is a monotone function on $t$. We will smoothly push the curve to the opposite side to make this integral constant.
@@ 5.1. Case $r=+\infty$
By Stokes theorem (see [26, Chapter IX, Theorems 5A and 7A])
@@ 8. Contactomorphisms
We continue to use Convention 2.9 assuming that the ambient manifold is equipped with some Riemannian metric.
We equip the set of all locally bi-Lipschitz self-homeomorphisms of the manifold with the compact-open topology. If the manifold is compact, this topology is induced by the standard metric
$\mathrm{dist}(f,g)=\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p)).$
Lemma 8.1.
The metric space of all $C$-bi-Lipschitz self-homeomorphisms of a compact manifold $M$ is complete for any $C\geq 1.$
Proof.
Pick a Cauchy sequence $f_{n}.$ It converges to a surjective continuous map $f.$ For any distinct $p,q\in M$
$\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(p,q)}=\lim\limits_{n\to\infty}\frac{\mathrm{dist}(f_{n}(p),f_{n}(q))}{\mathrm{dist}(p,q)}.$
Since $f_{n}$ are $C$-bi-Lipschitz, $f$ is a $C$-bi-Lipschitz homeomorphism. ∎
Remark 8.2.
Bi-Lipschitz homeomorphisms are given in local charts by Lipschitz functions. The set of all bounded Lipschitz functions on the open unit ball $B\subset\mathbb{R}^{n}$ coincides with the Sobolev space $W^{1,\infty}(B)$ (see [13] for details). It is natural to introduce a right-invariant metric on the space of bi-Lipschitz homeomorphisms which is topologically equivalent to the one induced by the Sobolev metric:
$\mathrm{dist}_{\mathrm{Lip}}(f,g)=\sup\limits_{p\in M}\mathrm{dist}(f(p),g(p))+\sup\limits_{p\neq q}\left|\ln\frac{\mathrm{dist}(f(p),f(q))}{\mathrm{dist}(g(p),g(q))}\right|.$
However this metric is left-discontinuous. So a continuous ambient isotopy can become discontinuous after multiplication from the left which is not convenient.
Definition 8.3.
A locally bi-Lipschitz self-homeomorphism of a smooth contact 3-manifold is called contact if it preserves the class of compact Lavrentiev Legendrian curves. We call such a homeomorphism a contactomorphism.
Recall that if $\beta$ is a 1-form on the smooth surface $S$ such that $d\beta$ is an area form, $\mathrm{pr}$ is the projection from $M=S\times\mathbb{R}$ to $S$ and $z$ is the projection $S\times\mathbb{R}\to\mathbb{R},$ then the kernel of the 1-form $dz+\mathrm{pr}^{*}\beta$ determines a contact structure on $M$. We equip $M$ with a product of a Riemannian metric on $S$ and the Euclidean metric on $\mathbb{R}$.
Lemma 8.4.
Let $h$ be a locally bi-Lipschitz homeomorphism of the smooth connected surface $S.$ Suppose that $\int\limits_{\gamma}\beta=\int\limits_{h(\gamma)}\beta$ for any closed rectifiable curve $\gamma.$ Let $p\in S.$ Then there is the unique contactomorphism $\widetilde{h}$ of $M=S\times\mathbb{R}$ such that $\widetilde{h}(p,z)=(h(p),z)$ for any $z\in\mathbb{R}$ and $\mathrm{pr}\circ\widetilde{h}=h\circ\mathrm{pr}.$
Proof.
Let $q\in S$ and $q\neq p.$ We connect the points $p$ and $q$ by a Lavrentiev arc $\gamma$ directed from $p$ to $q.$ By Lemmas 3.11 and 3.13 for any $z\in\mathbb{R}$ there exists the unique Legendrian curve having the endpoint $(q,z)$ and the projection $\gamma.$ The other endpoint of this curve is $(p,z+\int\limits_{\gamma}\beta).$ Since a contactomorphism must map Legendrian curves to Legendrian curves, we must have
The right side of this formula does not depend on $\gamma$ by our assumptions on $h$. So we can take this formula as the definition of $\widetilde{h}.$
$\widetilde{h}$ is locally bi-Lipschitz because the function $q\mapsto\int\limits_{\gamma}\beta-\int\limits_{h(\gamma)}\beta$ is locally Lipschitz. $\widetilde{h}$ maps Legendrian curves to Legendrian curves by construction. Therefore $\widetilde{h}$ is a contactomorphism. ∎
Example 8.5.
Let $\rho,\varphi,z$ be the cylindrical coordinates on $\mathbb{R}^{3}.$ Consider the contact form $dz+\rho^{2}d\varphi.$ A locally bi-Lipschitz homeomorphism of $xy$-plane preserves the integral of $\rho^{2}d\varphi$ on any closed rectifiable curve if and only if it preserves the area form $d(\rho^{2}d\varphi)=2dx\wedge dy.$ So the map
is a contactomorphism by the previous lemma. We note that the rays $\varphi=\mathrm{const},z=0$ are mapped to Legendrian curves whose projections to the plane $z=0$ are logarithmic spirals.
@@
The space of all $C$-bi-Lipschitz contactomorphisms is closed in the space of all $C$-bi-Lipschitz homeomorphisms for any $C\geq 1.$
@@
Suppose that a sequence of $C$-bi-Lipschitz contactomorphisms $f_{n}$ converges to a $C$-bi-Lipschitz homeomorphism $f$. Let $L$ be any compact Lavrentiev Legendrian curve. We need to prove that $f(L)$ is Legendrian. It is clear that $f(L)$ is a Lavrentiev curve.
Corollary 8.7.
The metric space of all $C$-bi-Lipschitz self-contactomorphisms of a compact contact manifold $M$ is complete for any $C\geq 1.$
@@
This is an immediate consequence of Lemma 8.1 and Proposition 8.6. ∎
Corollary 8.7 allows one to construct contactomorphisms as limits in $C^{0}$-topology.
@@
Let us prove that a smooth contactomorphism preserves the class of compact Legendrian Lavrentiev curves. Recall that Legendrian curves are characterized by the fact that the integral of the contact form on any their subarc is zero. Since the inverse image under the contactomorphism of the contact form is the contact form multiplied by some smooth nonzero function, the integral along any Legendrian arc remains to be zero after applying the contactomorphism. ∎
@@
We say that a family of maps from $X$ to $Y$ is locally uniformly (bi-)Lipschitz if for any compact $K\subset X$ there exists a number $C$ such that the restriction to $K$ of any map is $C$-(bi-)Lipschitz.
Definition 8.10.
A locally uniformly bi-Lipschitz ambient isotopy of a contact manifold is called contact if every homeomorphism in this isotopy is a contactomorphism.
Lemma 8.11.
Let $h_{t}$ be a locally uniformly bi-Lipschitz continuous family of homeomorphisms of a connected surface as in Lemma 8.4. Then the lift $\widetilde{h}_{t}$ constructed in that Lemma of this isotopy is a contact isotopy.
Proof.
For the proof it is only needed to check that the isotopy $\widetilde{h}_{t}$ is locally uniformly bi-Lipschitz. This is equivalent to the fact that the functions
$q\mapsto\int\limits_{\gamma}\beta-\int\limits_{h_{t}(\gamma)}\beta$
are locally uniformly Lipschitz. This is true because $h_{t}$ are uniformly Lipschitz. ∎
Figure 9. The bi-Lipschitz homeomorphisms from Example 8.12
@@
By Lemma 8.11 the following is a contact isotopy of the cylinder $\rho\leq 1$ in $\left(\mathbb{R}^{3},dz+\rho^{2}d\varphi\right)$:
$(\rho,\varphi,z)\mapsto\left\{\begin{aligned} \left(\rho,\varphi+\ln\rho,z-\frac{\rho^{2}}{2}+\frac{t^{2}}{2}\right),&\quad\rho\geq t,\\ \left(\rho,\varphi+\ln t,z\right),&\quad\rho<t,\end{aligned}\right.$
where $t\in[0;1].$ See Figure 9 for the images of polar rays under the $xy$-part of this isotopy.
@@
By Definition 8.10 the contactomorphisms $h_{t}$ are bi-Lipschitz with a common constant. Therefore the curves $h_{t}(L)$ are Lavrentiev with a common constant. So $\{h_{t}(L)\}$ is a Legendrian isotopy by Proposition 3.10. ∎